2ty/t^2+1 -2t(2-ln(t^2+1))y'=0​


1
Expert's answer
2021-05-07T14:39:09-0400

Given,

"(4+t^2)\dfrac{dy}{dt}+2ty=4t"


"\\[9pt] \dfrac{dy}{2-y}=\dfrac{2tdt}{4+t^2}"


Integrating Both the sides-


"-ln(2-y)=ln(t^2+4)+ln c_1

\\[9pt]

or c=\dfrac{2-y}{t^2}{4}\\[9pt]



So ,y=2-\dfrac{c}{t^2+4}=\dfrac{2t^2+8-c}{t^2+4}"